• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT
    Make New Entry, Make Followup Entry

    User name A. Camsonne

    Log entry time 02:31:14 on August 28, 2009

    Entry number 286428

    keyword=Owl shift run plan

    Also available at

    Run plan Aug 28th 2009
    From Hall A Wiki
    Jump to: navigation, search

    Current short term run plan

    * Owl shift ( August 28th 2009 )
    o complete Moller commissioning ( up to 1 AM )
    o Beam check Raster check ( set raster to 2mm x 2mm ) Check
    with spot and post in HALOG
    o HAPPEX detector Check ( Dustin, Rupesh ) ( ~ from 1 AM to 3 AM )
    + HAPPEX detector footprint alignement and Q2 measurement

    Move to Hydrogen target, check rates. Choose current reasonable for
    footprint determination ( few KHz of S0 triggers ) most likely around 1 uA.

    + Optics

    move Sieve in on both arms Increase current to have maximum T2 S0 up to 2
    KHz keeping current less than 30 uA. Keep an eye on the heater always
    keep 50 watt of overhead power.

    1 Sieve run with Hydrogen 1M events or 30 minutes whichever comes first

    Move to optics target, turn raster off, increase current up to 70 uA to
    get as close as possible to 2KHz of triggers

    Take on sieve run with Carbon optics ~ 1 Million events or 30 minutes
    whichever comes first.

    Move to Al dummy target Take runs to check position of Al superelastic
    until Dustin is happy. he can take an additional runs on Hydrogen if
    needed. Depending on the data he can make an access to move the detector.

    o BCM/BPM cross calibration and checks ( Kent / Luis ) Page
    Kent about 30 minutes before previous
    + BCM high current calibration of HAPPEX with UNSER
    + BPM calibration
    o Test of dithering coils ( Kent )

    This should take us up to day shift.

    Rupesh test will be done tomorrow during a down since he needs no beam
    maybe during low current Moller runs.

    A copy of this log entry has been emailed to: paschke@jlab.org,kkumar@jlab.org